OpenStudy (anonymous):

Find the volume of the composite solid if the diameter of the cone is 18.6 inches and the diameter of the cylinder is 22.5 inches. Round your answer to the nearest hundredth. There is a cone on top of a cylinder. The height of the cylinder is 12.07, and the side height, I think it's called the lateral height of the pyramid is 11.8 can you please help me find the volume of the composite solid if the diameter of the cone is 18.6 inches and the diameter of the cylinder is 22.5 inches. Round your answer to the nearest hundredth.

OpenStudy (campbell_st):

find the radius of the cylinder and cone radius= diameter/2 thats where you need to start the volume formulas are Cone \[V = \frac{1}{3}\times \pi \times r^2 \times h\] the cylinder \[V = \pi \times r^2 \times h\] in both cases h is the height of the cone and the cylinder